ne the molecular formulas to which the following em- c. CoC,Os (341.94 g/mol) d. n2. Derm to mulas and molar ade) and molar masses pertain. a. SNH (188.35 g/mol) h. NPCI, (347.64 g/mol) tem- red d. SN (184.32 g/mol) only carbon, hydrogen, and oxygen l for- 83. A compound that contains is 48.64% C and 8.16% H by mass, what is the empirica be mula of this substance? Themost common form of nylon (nylon-6) is 63.68% carbon. 12.38% nitrogen, 9.80% hydrogen, and 14.14% oxygen. Cal- 94, culate the empirical formula for nylon-6. 85. There are two binary compounds of mercury and oxygen. Heat- ing either of them results in the decomposition of the compound, with oxygen gas escaping into the atmosphere while leaving a residue of pure mercury. Heating 0.6498 g of one of the com- pounds leaves a residue of 0.6018 g. Heating 04172 g of the other compound results in a mass loss of 0.016 g. Determine the empirical formula of each compound. 86. A sample of urea contains 1.121 g N, 0.161 g H, 0.480 g C, and 0.640 g O. What is the empirical formula of urea? 87. A compound containing only sulfur and nitrogen is 6 6 S
Answer #1

82)

a)

Molar mass of SNH,

MM = 1*MM(S) + 1*MM(N) + 1*MM(H)

= 1*32.07 + 1*14.01 + 1*1.008

= 47.088 g/mol

Now we have:

Molar mass = 188.35 g/mol

Empirical formula mass = 47.088 g/mol

Multiplying factor = molar mass / empirical formula mass

= 188.35/47.088

= 4

Hence the molecular formula is : S4N4H4

b)

Molar mass of NPCl2,

MM = 1*MM(N) + 1*MM(P) + 2*MM(Cl)

= 1*14.01 + 1*30.97 + 2*35.45

= 115.88 g/mol

Now we have:

Molar mass = 347.64 g/mol

Empirical formula mass = 115.88 g/mol

Multiplying factor = molar mass / empirical formula mass

= 347.64/115.88

= 3

Hence the molecular formula is : N3P3Cl6

c)

Molar mass of CoC4O4,

MM = 1*MM(Co) + 4*MM(C) + 4*MM(O)

= 1*58.93 + 4*12.01 + 4*16.0

= 170.97 g/mol

Now we have:

Molar mass = 341.94 g/mol

Empirical formula mass = 170.97 g/mol

Multiplying factor = molar mass / empirical formula mass

= 341.94/170.97

= 2

Hence the molecular formula is : Co2C8O8

d)

Molar mass of SN,

MM = 1*MM(S) + 1*MM(N)

= 1*32.07 + 1*14.01

= 46.08 g/mol

Now we have:

Molar mass = 184.32 g/mol

Empirical formula mass = 46.08 g/mol

Multiplying factor = molar mass / empirical formula mass

= 184.32/46.08

= 4

Hence the molecular formula is : S4N4

I am allowed to answer only 1 question at a time